OpenStudy (anonymous):

consider the function h = 100sin {(pi/240)t} + 50 For what values of t is h>100 where 0

OpenStudy (phi):

you want \[ 100 \sin\left(\frac{\pi\ t}{240}\right) + 50 > 100 \] for 0< t < 350. that means the angle is between 0 and 1.46\(\pi\) radians a sketch of the sin function shows it is positive between 0 and π radians The next step is solve for sin() to get sin() > value then take the inverse sin of both sides. Look at the sine function to get an idea of the angles that satisfy the relation